OpenStudy (precal):

Disk Method: The bounded region formed by y=x^(1/2) , y = 2, and the y-axis is rotated about the line y = 2. Find the volume of the solid.

OpenStudy (precal):

I know this is dx

OpenStudy (precal):

\[\pi \int\limits_{0}^{4}\left( 2-\sqrt{x} \right)^2 dx\]

OpenStudy (precal):

Is this the correct set up? @sourwing

OpenStudy (phi):

yes

OpenStudy (precal):

I got (8/3) pi

OpenStudy (precal):

Thanks phi